### v1

# Agentic Visualization: Extracting Agent-based Design Patterns from Visualization Systems

Abstract
Autonomous agents powered by Large Language Models are transforming AI, creating an imperative for the visualization field to embrace agentic frameworks. However, our field’s focus on a human in the sensemaking loop raises critical questions about autonomy, delegation, and coordination for such agentic visualization that preserve human agency while amplifying analytical capabilities. This paper addresses these questions by reinterpreting existing visualization systems with semi-automated or fully automatic AI components through an agentic lens. Based on this analysis, we extract a collection of design patterns for agentic visualization, including agentic roles, communication and coordination. These patterns provide a foundation for future agentic visualization systems that effectively harness AI agents while maintaining human insight and control.
journal: IEEE Computer Graphics & Applicationsfootnotetext: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
[…]
Autonomous agents—software entities that perceive their environment and take goal-directed actions—have been a cornerstone of artificial intelligence research for decades [22]. The emergence of Large Language Models (LLMs) has dramatically accelerated this paradigm, enabling general-purpose agentic AI systems capable of sophisticated planning, reasoning, and autonomous action in complex domains [20]. Given how previous waves of AI innovation have transformed visualization, this agentic turn will inevitably reshape our field. Yet visualization differs fundamentally from other AI application domains through its emphasis on human-centered reasoning, support for open-ended sensemaking, and placement of human analysts at the center of the analytical process. This creates an inherent tension between autonomous agent capabilities and visualization’s human-centered focus, raising questions about agency, delegation, coordination, and responsibility [6]. How can we design visualization systems that leverage agentic capabilities while preserving human insight and analytical control? What visualization tasks can we confidently delegate to agents?
To address these questions, we introduce the concept of agentic visualization (AV): interactive visual analysis systems that incorporate autonomous AI agents while maintaining human agency in the analytical process. Rather than starting from scratch, we analyze existing visualization systems with automatic or semi-automatic capabilities—sometimes described as mixed-initiative systems [7]—reinterpreting them as early forms of agentic visualization. By examining these systems through an agentic lens, we identify recurring structures that effectively coordinate automated components and human analysts. We formalize these structures as design patterns for agentic visualization that balance automation and analytical control. Design patterns, originating in architecture [1] and software engineering [5], provide a structured vocabulary for describing reusable solutions to common challenges. This approach aligns with human-centered AI principles that emphasize human values, interpretability, and agency in AI-augmented systems [2, 13].
We validate our proposed design patterns by identifying their presence across diverse published visualization systems. Each pattern is structured with a descriptive name, the problem it addresses, its solution structure, implementation examples, and associated tradeoffs. To demonstrate practical application, we present a novel example scenario that combines multiple patterns to create a dynamic visualization dashboard surveying an active scientific field where agents with different roles, communication, and coordination mechanisms work together to update progressively.
This paper makes the following contributions: (1) The concept of agentic visualization as a framework for designing and analyzing systems that incorporate autonomous agents while preserving human-centered analytical processes; (2) a collection of design patterns for agentic visualization extracted from a systematic analysis of prior literature on automated and semi-automated visualization systems; and (3) a demonstration showing how these patterns can be applied to a novel analytical scenario, including their composition into higher-order solutions. {mdframed}[backgroundcolor=blue!10, frametitle=
Sidebar: Agents and AI
, frametitlerule=true, frametitlebackgroundcolor=bgcolor]
Autonomous agents—software entities that perceive their environment and take goal-directed actions—have been a cornerstone of artificial intelligence research since the field’s inception in the 1950s [22]. The early reactive agents of the 1990s emphasized real-time interaction over complex reasoning, exemplified by Rodney Brooks’ subsumption architecture [17], followed by multi-agent systems and cognitive architectures that aimed to model distributed problem-solving and human cognition respectively [18].
The current resurgence of agent technology, powered by Large Language Models (LLMs), has dramatically accelerated this paradigm, enabling general-purpose agentic AI systems capable of sophisticated planning, reasoning, and autonomous action in complex domains [20]. This represents a significant shift in capabilities but confronts persistent challenges that historical agent systems also faced: poor generalization across domains, scalability issues with complex tasks, inadequate coordination mechanisms, brittle performance in unexpected situations, and ethical concerns. These limitations highlight an inherent tension in agent design that was famously debated by Ben Shneiderman and Pattie Maes at IUI ’97 and CHI ’97, where Shneiderman advocated for direct manipulation interfaces that “affords the user control and predictability,” while Maes argued for delegation to software agents that could learn user preferences and act on their behalf [21]. This debate continues to shape the balance between computational agency and human control in interactive systems.
Addressing these challenges requires more than just technical improvements to agent capabilities. As Shah and White argue, successful agent systems need to generate sufficient value for users, provide adaptable personalization, establish trustworthiness, achieve social acceptability, and develop standardization protocols. Aridor and Lange’s work on design patterns for agent systems provides a useful foundation for capturing common solutions to recurring problems in agent design [16]. Their patterns are conceptually divided into three classes:
  • Traveling patterns: Agents moving and navigating in distributed environments;
  • Task patterns: Accomplishing agent objectives and distributing workloads; and
  • Interaction patterns: Governing communication and coordination between agents.
Recent work has extended this approach to foundation model-based agents, with researchers developing catalogs of architectural patterns that address challenges such as hallucinations, explainability, and accountability [19].
RELATED WORK
The notion of agents lies on a spectrum of varying degrees of their inherent agency over their own actions.
On one end of that spectrum, agents are explicitly called upon by the user and given the autonomy to perform tasks under changing and unanticipated conditions without a user’s oversight. This “autonomous systems” approach is embodied in early visualization research dating back to the 1990s. It can be found in concepts like Smart Particles that are sent out into the data space, travel through it, and in their wake leave behind visual elements that make the data space visible as the swarm of particles wanders across it [10]. Another example is the idea of Design Agents with different strategies and abilities and which by interacting with each other, yield visualizations as an emergent joint outcome that automatically adjusts to underlying data changes as the agents react to them [8].
The next step on that spectrum are agents that continuously work alongside the users without being explicitly triggered. They silently observe the users’ actions and offer unobtrusive help, possibly even extrapolating their needs and speculatively carrying out anticipated tasks ahead of time. This echoes ideas such as Mark Weiser’s Calm Computing, sometimes likened to “a raft of invisible servants” [11]. In visualization, we can find this idea in agent-based systems that run alongside or inside a visualization system and gently aid users, for example, in finding multi-view setups that maximize visibility and relevance of the shown data [9] or by suggesting meaningful next steps in visual analyses [14].
On the other end of the spectrum, these independent agents are proactive themselves and may even be permitted to interrupt the user if needed. Visualization research has long explored such mixed-initiative interactions [7] where computational systems and human analysts divide the labor and collaboratively build understanding. In progressive visualizations, for example, agents may monitor in-progress views for changes, alerting the users in case any relevant changes occur—in particular in off-screen views [29]. This frees the human users from monitoring and allows them to explore the data freely without fearing to miss out on anything happening elsewhere. Similarly, independent processes can be used to check a visualization design for construction errors and readability issues (visualization linters), or check an ongoing visual analysis for potential biases—again alerting users if any issues are detected.
AGENTIC VISUALIZATION
We define agentic visualization as the integration of autonomous AI agents with interactive visual analysis systems while balancing computational autonomy with human analytical control. This approach enables agents to independently perform complex analytical tasks while preserving the human analyst’s ability to guide the analytical process, interpret results, and maintain decision-making authority.
Agentic visualization extends and transforms several existing paradigms in visual analytics. Building on mixed-initiative interaction [7] and human-AI collaboration [2], agentic visualization specifically emphasizes agent autonomy beyond reactive assistance. Beyond merely responding to explicit user actions, agentic visualization allows for proactively identifying opportunities for analysis, suggesting approaches, or even pursuing analytical threads independently.
This shift challenges traditional interaction models where users issue commands and wait for responses. Instead, agentic visualization establishes a more fluid relationship where both human and agent can initiate analytical activities. This raises important questions about idle behavior: should an agent remain passive until directed (cold start), or should it proactively explore data and generate insights even without explicit user direction? The answer likely depends on context, task complexity, and user preferences.
Agentic visualization also leverages progressive visual analytics [15], where systems provide incremental visual updates during computationally intensive processes. However, while progressive analytics primarily focuses on making computation visible, agentic visualization extends this by enabling agents to make autonomous decisions about which computational paths to pursue based on intermediate results. This way, agents can prioritize promising analytical directions, abandon unproductive paths, and adapt strategies without continuous human supervision.
But this evolution raises a critical question: how do we design agentic visualization systems that effectively balance agent autonomy with human control? This is particularly important in a field like visualization, which emphasizes the role of the human analyst in the interactive loop. Accordingly, the visualization field currently lacks systematic guidelines for determining appropriate levels of agent initiative, establishing effective communication between agents and analysts, and maintaining analytical transparency when agents operate with increasing independence.
To address these challenges, we look to past literature but reframe it through an agentic lens. By examining existing visualization systems with automatic or semi-automatic capabilities, we can identify patterns of effective human-agent collaboration that have emerged organically. These patterns, once formalized, can provide a foundation for designing the next generation of agentic visualization systems.
PATTERNS FOR AGENTIC VISUALIZATION
Agentic Visualization Patterns.
Figure 1: Overview of agentic visualization patterns. Agent role patterns running their computations throughout the data analysis workflow and interacting with various human roles, based on Crisan et al. [4].
To systematically identify patterns for agentic visualization, we conducted a comprehensive analysis of visualization literature where autonomous or semi-autonomous components collaborate with human analysts. Our methodology focused on extracting recurring design solutions that effectively balance computational agency with human control in visual analytics systems.
Our review focused on prior work that met four specific inclusion criteria: (1) involves direct human user interaction; (2) has an emphasis on data and data analysis; (3) includes interactive visualization components; and (4) contains autonomous or semi-autonomous computational elements. We did not limit our analysis to systems explicitly labeled as “agent-based,” recognizing that many visualization systems incorporate autonomous behaviors without using this terminology. Instead, we focused on functional characteristics that exemplify human-agent collaboration in visual analytics contexts. Furthermore, we restricted our search to academic publications for this survey.
From this analysis, we identified patterns that appeared in at least two distinct systems, suggesting their broader applicability. We also selected patterns that, while perhaps less common in the literature, show particular promise for agentic visualization research.
Following long-standing design pattern praxis [1, 5], each pattern in our catalog has a descriptive title capturing the pattern’s essence; the specific problem the pattern addresses; the reusable solution that resolves the problem; and the consequences of implementing the pattern, including benefits, limitations, and tradeoffs. Our analysis revealed three distinct categories of design patterns for agentic visualization:
  1. 1.
    Agent Role Patterns: These patterns define the fundamental roles and responsibilities agents can assume within a visualization system. They establish the possible scopes of agent autonomy and the division of analytical labor between human and computational actors.
  2. 2.
    Communication Patterns: These patterns structure the information exchange between agents and human analysts. They address how agents present findings, receive feedback, and maintain transparency about their operations and reasoning processes.
  3. 3.
    Coordination Patterns: These patterns govern the temporal and workflow dynamics between agent and analyst activities. They establish when and how agents act, how concurrent human-agent actions are synchronized, and how analytical control shifts between participants.
Table 1 presents our proposed catalog of agentic visualization patterns organized by category. This catalog is not exhaustive but provides a foundation for understanding and designing effective agentic visualization systems. In the following sections, we elaborate on each pattern category and detail the specific patterns within them, providing concrete examples from existing systems and discussing their implications for future research and development in agentic visualization.
AV Category Pattern Examples
Role Patterns Forager [27, 38]
Analyst [36, 40]
Chart Creator [32, 30]
Storyteller [26, 35]
Communication Patterns Insight Timeline [23, 25]
Progress Indicator [23, 29]
Provenance Log [24, 33]
Coordination Patterns Scouting [10, 25]
Swarming [34, 31]
Monitoring [23, 29]
Consolidating [39, 35]
Table 1: Overview of our agentic visualization patterns. The identified patterns and their corresponding categories. For each pattern, we give several examples from existing visualization systems that showcase the pattern.
{mdframed}
[backgroundcolor=blue!10, frametitle=
Sidebar: Visualization and Agents
, frametitlerule=true, frametitlebackgroundcolor=bgcolor]
While agents have not been a major theme in the scientific community for visualization so far, there have been several instances of autonomous and semi-autonomous entities in visualization systems in the literature. Here we review these prior systems and their autonomous components.
  • InsightsFeed [23] is a progressive visual analytics system that generates and continuously refines an analysis of social media posts based on keywords, sentiment, and multidimensional embedding.
  • VisTrails [24] manages visualization workflows and their provenance metadata, enabling reproducibility, comparison of visualization pipelines, and systematic tracking of the scientific discovery process.
  • DataSite [25] is a proactive visual analytics system that autonomously executes background computations to identify salient data features and surfaces these findings as notifications in a feed timeline.
  • DataWeaver [26] is a data story authoring system that supports vis-to-text and text-to-vis compositions through statistics, custom algorithms, and LLMs.
  • SurfaceMapper [27] identifies relevant surfaces in 3D scientific data through multiple complementary agents that perform independent data interpretations.
  • HaLLMark [28] is a writing assistance system that autonomously captures and visualizes the complete provenance of writer-LLM interactions.
  • ProReveal [29] is a progressive visual analytics system that autonomously monitors intermediate analytical findings and activates safeguards when uncertain knowledge requires verification.
  • AVA [30] is a visualization agent framework that uses multimodal LLMs to autonomously interpret visual outputs and make visualization parameter decisions.
  • Medley [32] is a mixed-initiative system that recommends visualizations and widgets for dashboard composition, automatically providing interaction techniques.
  • StoryFacets [33] is a collaborative analysis and storytelling system that autonomously maintains visual provenance across multiple linked views.
  • Data Director [35] is a multi-agent system where LLM-powered agents generate insights, visualizations, and text to integrate them as a story into animated data videos.
  • Snowy [36] is a natural language interface that generates contextual utterance recommendations by combining data interestingness metrics and language pragmatics to guide analytic workflows.
  • DynaVis [37] is a multimodal system that combines a natural language interface with WIMP elements to generate standard charts and provide automatically synthesized widgets for visualization editing.
  • In-formation flocking [38] is a visualization approach that leverages a decentralized multi-agent system to recognize dynamic patterns and clusters in time-varying datasets, with local and global views.
  • InsightLens [39] is an LLM-agent-based interactive system that automatically captures and manages insights from users’ conversational data analysis workflow.
  • Voyager [40] generates galleries of visualization recommendations based on statistical and perceptual measures, enabling analysts to explore data without manual specification of variables or encodings.
AV ROLE PATTERNS
We identified four main agent roles for visualization.
[…]
Pattern #1: The Forager pattern describe agents that look for data of interest according to parameters provided by the human. This pattern is associated with the challenges of processing and interpreting large datasets while keeping an overview of the dataset and adjusting to the domain context. Forager agents often work in groups, yet each Forager can either apply custom rules or variations of the given parameters to work efficiently in parallel and report back the results.
[…]
*[regular] Examples: In SurfaceMapper [27], multiple agents with complementary expertise cooperate to identify data of interest in large volumes of scientific data. The agents report their findings to the human user and exchange them to revise their hypotheses. The in-formation flocking [38] technique involves multiple Forager agents that follow a set of behavior rules with the goal of highlighting potential patterns of interest in time-varying datasets. Each agent focuses on specific data neighborhoods and calculates similarity metrics to detect temporal changes. Working with Forager agents reduces the manual work of data processing and interpretation by collecting samples of interest and preliminary insights that the human can later compare and synthesize. Due to their flexibility, these agents can adjust to different domains. However, users still have to configure the agents according to their needs.
[…]
Pattern #2: An Analyst agent generates and guides human users in making sense of a dataset. This pattern is associated with the challenges of exploring any data without specific goals to steer the process. To assist people who may struggle to choose a starting point for their exploration (i.e., the cold start problem), Analyst agents recommend multiple ways to start. The agent analyzes the dataset and prepares an initial set of recommendations. As soon as the person starts to interact, the agent starts learning from those actions to make follow-up recommendations. The Analyst encourages breadth-oriented exploration to increase data and analytical coverage.
[…]
*[regular] Examples: Voyager [40] provides visualizations of all dataset variables to give an overview of the data space so the human user can choose what variables to focus on. In contrast, Snowy [36] generates multiple natural language inquiries based on patterns detected automatically in the dataset. The support of a computational Analyst allows users to have fewer concerns about leaving areas unexplored or missing key patterns. However, dealing constantly with updated recommendations may be overwhelming and distract the person from focusing on a specific data sample. Future work on Analyst agents should look for ways to balance these tradeoffs, adapting to the person’s skills.
[…]
Pattern #3: Chart Creator agents assist or co-create visualizations with the human user. This pattern is associated with authoring scenarios, where the goal is to create and refine one or multiple visualizations. This involves a series of decisions, such as choosing an appropriate visual encoding and adjusting visualization parameters. When creating composite visualizations, such as a dashboard, choosing what visualizations to combine is also necessary. These choices are often made through multiple trials and repetitive actions. They require visualization literacy as well as technical knowledge, depending on the tool. Chart Creator agents help reduce the workload by helping the authoring process.
[…]
*[regular] Examples: In DynaVis [37], the agent creates charts on demand and generates editing widgets when the person asks to change a visualization parameter. Medley [32] provides visualizations and widgets grouped by analytical intent that users can select from when authoring a visualization dashboard. The agent takes care of configuring standard interactions. One challenge of this pattern is that interacting with the agent is often the only way to create visualizations, potentially adding an extra layer of complexity for simple changes. Moreover, the automatically created elements are constrained to specific visualizations and widgets. Still, co-creating with an agent can reduce the need for repetitive tasks and advanced knowledge.
[…]
Pattern #4: Storyteller patterns provide an agent who creates and updates data-driven stories. This pattern addresses the problem of generating and composing a cohesive story that brings together text, insights, and visual elements. Given that these processes require multiple skills regarding data analysis, text writing, and visualization authoring, the agent complements the skills of the human user by taking over one or multiple tasks. These agents help in the storytelling workflow by creating specific story elements.
[…]
*[regular] Examples: DataWeaver [26] generates narrative text through dedicated algorithms and LLMs and integrates them into the main story. Data Director [35] provides an agent for designing and synchronizing visual animations for a data video. Designating storytelling tasks to agents helps people focus on specific story elements by delegating the generation or refinement of others. This can speed up the process and reduce barriers for people who may only have expertise in text writing, data wrangling, or visuals creation. The corresponding tradeoff is that the agent’s output needs to be checked for accuracy and appropriateness, which creates a new task for the human user and may represent a challenge if the person does not have the corresponding expertise to evaluate it.
AV COMMUNICATION PATTERNS
Agents deployed for addressing visualization-specific tasks—from analysis to storytelling—should also communicate and present their findings to the user. Here we discuss these communication patterns through which agents and users share task allocation, progress, and completion.
[…]
Pattern #5: The Insight Timeline pattern deals with situations where multiple agents, both humans and AI-powered ones, need a way to synchronize and share findings while continuing to work autonomously on their own tasks. The solution is to provide a common channel for agents and humans to share new insights and findings collected during the data analysis process. An Insight Timeline serves as a channel where both ongoing and explored tasks can be shared to (i) declare the current focus of exploration within the dataset to avoid redundancy, (ii) report interesting findings, and (iii) steer the analysis direction.
[…]
*[regular] Examples: InsightsFeed [23] explores large textual datasets and progressively updates the results in an Insight Timeline so that the user can steer their analysis direction based on the results shared by the system. Similarly, DataSite [25] adopts the UI principles from InsightsFeed to propagate the results of background computation to the user without interrupting the cognitive process of the user. Users can also rank, sort, and filter the insights based on their needs.
[…]
Pattern #6: The Progress Indicator pattern supports communication about which tasks the agents have undertaken and how far along each task is. This pattern helps human analysts interpret intermediate results, assess uncertainty in the current analysis, and identify areas of interest.
[…]
*[regular] Examples: InsightsFeed [23] shows the progress and the aliveness by using a blinking rectangle next to a progress bar to highlight the current chunk of data being processed by the system. The system also uses a quality/stability metric to indicate the quality or uncertainty of the current state. They visualize the quality metric as a small line chart on top of each visualization to indicate if the quality of the computation is increasing, decreasing, or stable for the data behind the visualizations. ProReveal [29] shows the progression of both ongoing and completed visualizations to the analyst using a progress ring, which is hidden once the visualization is completely rendered with all the data. In both InsightsFeed and ProReveal, users have the option to interact with this process through pause, resume, and remove buttons to control the progression and explore the data in its current state. In addition to task progression, quality, and uncertainty, the progress indicators can also be used for other measures such as time spent vs. total time left, the amount of data processed, and the speed of the task performance.
[…]
Pattern #7: The complex and collaborative data analysis process makes it difficult to track what steps were performed, by whom and when. To ensure reproducibility, verification, and replication of analysis steps, the Provenance Log pattern provides a common audit log of sources, transformations, and filters applied to the data by both human and AI agents.
[…]
*[regular] Examples: VisTrails [24] captures provenance for visualization by tracking all the changes that users make to the visualization pipeline in an exploration task. Users can interact with the provenance information—presented as a directed graph—to reproduce a specific state and understand the process history. For collaborative exploration, VisTrails allows users to integrate other analysts’ workflows afterward in their view to analyze the provenance information. StoryFacets [33] presents a simpler view for displaying provenance from multiple collaborators on the same analysis session by creating multiple linked formats called facets. HaLLMark [28] uses summary statistics to encode the percentage of tasks performed by human and AI in a collaborative writing process.
AV COORDINATION PATTERNS
Agentic visualization workflows often employ multiple agents. Depending on the nature of the task, agents may work towards a shared goal, operate independently, or wait for others to complete their tasks before proceeding. Coordination is essential to ensure that agents and human analysts can schedule and execute their tasks effectively. Here we discuss different coordination patterns that describe how agents and human analysts coordinate during analysis.
[…]
Pattern #8: At the beginning of the analysis process, it is often unclear which tasks should be prioritized to perform the analysis effectively, especially in large datasets. Scouting can serve as an initial coordination pattern, where multiple autonomous agents independently explore the dataset to identify potentially interesting patterns, anomalies, or other areas of interest. Human analysts can provide a single or multiple goals, and the agents can operate with this knowledge and adapt to the analysts’ feedback.
[…]
*[regular] Examples: Mix&Match [10] releases smart particles (i.e., agents) into the dataset to highlight features of interest. The particles operate on rules, have a finite lifetime and scout for information in the dataset, leaving behind visual effects for the users. Users can also modify particle behavior by updating the rules to produce different effects. DataSite [25] proactively runs background computations on dataset to support human analysts during the different stages of the visual data exploration process. The analyst still maintains agency over the exploration and is actively supported by the system to generate insights, creating a balanced dialogue between the analyst and the system (or agents) and utilizing their respective strengths.
[…]
Pattern #9: In some cases, it might not be enough to employ autonomous agents on independent tasks with different goals, but multiple agents should be deployed with a single task and a shared goal. Swarming agents can address this issue by dividing the problem into smaller components that each agent in the swarm can autonomously work on and coordinate.
[…]
*[regular] Examples: To solve a clustering problem, Saka et al. [34] designed a simultaneous clustering and visualization algorithm that uses an ongoing swarm of agents to coordinate the tasks with each other. They use a flock-based approach for data clustering where each agent receives a position in a 2D or 3D visualization panel. Similar agents come together to form swarms to visualize dynamic clustering where data is collected over time. Vande Moere [31] proposed visualization agents that determine their own visual properties through continuous interactions with other agents. Their iterative organization creates visual effects based on data properties that both agents and users can perceive. These agents use the concept of Swarming and can identify other agents in their vicinity, read their data and move towards or away from them based on their data properties. Using a stock dataset, where agents represent companies, Moere describe how agents flock together or apart based on the similarity of their price change behavior over time, leading to different patterns such as short-term clustering and long-term zoning patterns.
[…]
Pattern #10: Monitoring enables agents launched during the analysis process to track temporally streaming data and notify the analyst of any changes, without overwhelming the human analyst.
[…]
*[regular] Examples: InsightsFeed [23] enables analysts to keep track of evolving results during progressive data analysis. In addition to the visual exploration of incrementally processed data, users can also monitor the progress of the background computation and see how far along it is. DataSite [25] also uses background processes to support the human analysts in visual data exploration and help them actively monitor the results of these processes using a feed timeline. In ProReveal [29], analysts can start the exploration and employ agents with different analytical support to guard the analysis process as a large dataset is progressively explored. It uses a web-based user interface where human analysts can explore the dataset, employ the guards, and continue exploring while Monitoring the guards’ progress. They are notified when a change occurs and the condition is modified.
[…]
Pattern #11: As multiple agents perform various tasks throughout the analysis process, Consolidating the outcomes becomes essential to help both humans analysts and agents to build upon these findings. Based on the consolidated results, further exploration, summarization, or sensemaking of the data can be done. For example, after Consolidating the findings from Scouting the dataset, human analysts may wish to deepen their exploration in specific areas of interest. This is especially useful during sensemaking, where human analysts can interpret and communicate important findings from consolidated information.
[…]
*[regular] Examples: InsightLens [39] uses an LLM-agent-based framework that automatically records and organizes insights during conversational data analysis with a human and an LLM agent. Based on the outcome of the analysis, the user can monitor and navigate the conversation using visualizations and receive guidance on further data exploration without interrupting their workflow. DataDirector [35] uses multiple agents to automate the creation of animated videos from data. LLM-based agents autonomously interpret the data, analyze it and further spawns design agent roles to create final data videos. In this case, tasks are broken down into sub-tasks at each analysis step, and the results from these tasks serve as input to the next step. In the final step, the main controller agent consolidates the decisions made by the agents to generate a data video for the human analyst.
APPLYING AV PATTERNS
To demonstrate the practical application of our proposed AV patterns, we present an example scenario of a Medical Research Monitoring system that addresses the growing challenge of keeping pace with rapidly expanding medical literature. Consider a research team in a medical domain, where thousands of new publications emerge continuously across various sources that may be relevant to the field of study or a certain treatment or medication. The volume and velocity of this research make comprehensive manual monitoring impossible, yet staying current is critical. Traditionally, extensive manual efforts are required to identify key papers and extract significant information in order to visualize the literature landscape and synthesize the latest study results. This complex scenario leads to several interconnected challenges:
[…]
Challenge 1: Query formulation and seeding. Significant domain expertise is required to create comprehensive search strategies that capture relevant publications without overwhelming noise. The Forager pattern directly addresses this challenge by deploying multiple autonomous agents that can explore different search strategies simultaneously. Each Forager agent can be configured with complementary expertise, and these agents can autonomously refine their search parameters based on initial results. Additionally, the Analyst pattern can assist researchers in formulating initial queries by analyzing existing literature collections to suggest relevant search terms and identify potential gaps in current search strategies.
[…]
Challenge 2: Scale and filtering. The challenge of processing the overwhelming volume of publications that appear daily requires efficient pre-filtering mechanisms to identify potentially relevant studies while avoiding critical omissions that could impact research completeness. The Scouting coordination pattern provides a solution for handling the massive scale of incoming publications. Multiple scouting agents can be deployed to independently explore different publication streams, with each applying domain-specific filtering criteria while maintaining awareness of overall research priorities. This approach, combined with Monitoring, enables researchers to track the progress of filtering operations across multiple databases simultaneously. The Progress Indicator communication pattern keeps researchers informed about processing status without overwhelming them with intermediate results, allowing them to focus on high-priority publications while background agents handle the volume.
[…]
Challenge 3: Concept extraction and validation. The task of automatically identifying and extracting key findings, methodologies, and outcomes from diverse publication formats while ensuring accuracy and completeness across varied writing styles and reporting standards presents significant complexity. The Analyst pattern excels at automated concept extraction, using natural language processing and domain-specific knowledge to identify key findings across diverse publication formats. These agents can generate structured summaries of extracted concepts and present them through the Insight Timeline communication pattern. Additionally, Chart Creator agents can generate domain-specific visualizations such as temporal trends, citation networks, and medication-specific charts that highlight key elements and emerging topics. The Provenance Log pattern ensures that all extracted concepts maintain links to their source publications, enabling researchers to verify accuracy and trace findings back to original sources for validation.
[…]
Challenge 4: Research cluster identification. The task of discovering complex relationships between studies—including methodological similarities, conflicting results, and emerging research directions—requires sophisticated analysis of both explicit citations and implicit conceptual connections. The Swarming coordination pattern is particularly well-suited for discovering complex relationships between studies. Multiple agents can work collaboratively to identify clusters based on different similarity metrics—methodological approaches, research populations, outcome measures, or citation networks. Each agent in the swarm contributes its specialized perspective while coordinating with others to build an understanding of research relationships. The Insight Timeline pattern facilitates sharing of discovered clusters among team members, enabling collaborative validation and refinement of identified data patterns.
[…]
Challenge 5: Knowledge synthesis and gap analysis. The complex process of integrating findings across studies to maintain a coherent overview of the research landscape and systematically identify opportunities for future research while managing contradictory evidence and varying quality levels presents ongoing challenges. The Storyteller pattern addresses the complex task of synthesizing knowledge across multiple studies. These agents can generate coherent narratives that integrate findings from related research clusters, identify contradictory results, and highlight emerging trends. The Consolidating coordination pattern ensures that insights from multiple agents are systematically integrated into comprehensive research overviews. Combined with Provenance Log, researchers can trace the synthesis process and understand how specific conclusions were derived from the underlying evidence.
[…]
Challenge 6: Temporal monitoring and adaptation. The ongoing need for continuously updating the monitoring system as new research emerges and research priorities evolve requires adaptive strategies that can respond to shifting focus areas and changing research methodologies. The Monitoring coordination pattern provides continuous oversight of the evolving research landscape. Monitoring agents track changes in publication patterns, emerging topics, and shifting research priorities, alerting researchers when significant developments occur. The Progress Indicator pattern communicates the evolution of research areas over time, while the Insight Timeline serves as a living record of research developments. This temporal awareness enables the system to adapt its focus areas dynamically, ensuring that monitoring strategies remain aligned with current research needs and emerging opportunities.
RESEARCH VISION AND OUTLOOK
Agentic visualization systems must navigate the relationship between computational autonomy and human analytical control. As Shneiderman [13] points out, human versus AI control is not a zero-sum game, but can be thought of as a 2D design space where human and computer control are independent axes. This framing reveals opportunities where both human and AI agents can retain high degrees of control.
As AI systems continue to evolve, we will most likely witness the rise of agents throughout the visual analysis process. Due to their significant computational power, agents can address issues with scalability, reducing the time from data to insights. LLM-based agents have already taken on collaborator roles in fields like programming, acting as pair programmers, code reviewers, and automated testers. A similar collaborative role could extend to visual analytics, where agents support human analysts by relieving them of cognitively demanding or tedious tasks.
Consider scenarios where AI agents proactively provide background and contextual information for insights, make suggestions for analytical directions, and provide careful provenance and source information for all decisions while leaving ultimate decision-making and sensemaking to the human analyst. An Analyst agent might autonomously identify statistical outliers in a dataset, gather relevant contextual information about those outliers, and present this analysis with full methodological transparency; a Chart Creator might suggest one or several charts suitable for representing the findings; and a Storyteller might suggest a comprehensive storyline for presenting the high-level insights. However, in the end, the human analyst retains control over interpreting significance and choosing next steps. With communication approaches such as Insight Timeline, agents and human analysts will be able to interact seamlessly through natural and visual language alike.
Practically speaking, we suggest deploying agentic visualization for the lower parts of the sensemaking loop [12]—information foraging tasks at lower abstraction levels that involve data processing, pattern detection, and routine analytical operations. The higher, more cognitively demanding levels can then be reserved for human analysts: forming hypotheses, making contextual judgments, and drawing conclusions that require domain expertise and ethical reasoning.
The shift toward agentic visualization raises questions about analytical responsibility, interpretability, and the preservation of human values that have long defined our field. When an AI agent discovers a pattern that influences policy decisions, questions of responsibility become complex. How do we maintain transparency when agents make autonomous analytical choices? What decisions should agents be allowed to make? How do we ensure that increased automation enhances rather than undermines human analytical reasoning? These concerns extend beyond traditional AI ethics frameworks.
The patterns presented in this work may be helpful to address these issues. For example, Provenance Log can maintain comprehensive logs of data sources and transformations applied to the data by both agents and users. Such provenance trails could ensure transparency and reproducibility of the analysis process. Additionally, quality and uncertainty metrics could be employed [29] to indicate the quality of the results from agents. Still, ultimately, domain experts are required to validate the quality of the results.
Visualization has always emphasized human insight and interpretability as core values; put differently, the ability to “use vision to think” [3] is fundamental to our area. Accordingly, agentic systems must preserve these values while expanding analytical capabilities. This requires agents that can explain their analytical choices, expose their reasoning processes, and maintain clear audit trails of autonomous actions. As analysts, we will still need to be able to think through the evidence provided to our eyes. The patterns and principles outlined in this paper aim to guide us in achieving this balance.
Acknowledgments
This work was supported partly by Villum Investigator grant VL-54492 by Villum Fonden. Any opinions, findings, and conclusions expressed in this material are those of the authors and do not necessarily reflect the views of the funding agency.