### v3

# Agentic Visualization: Extracting Agent-based Design Patterns from Visualization Systems

Abstract
Autonomous agents powered by Large Language Models are transforming AI, creating an imperative for the visualization area. However, our field’s focus on a human in the sensemaking loop raises critical questions about autonomy, delegation, and coordination for such agentic visualization that preserve human agency while amplifying analytical capabilities. This paper addresses these questions by reinterpreting existing visualization systems with semi-automated or fully automatic AI components through an agentic lens. Based on this analysis, we extract a collection of design patterns for agentic visualization, including agentic roles, communication, and coordination. These patterns provide a foundation for future agentic visualization systems that effectively harness AI agents while maintaining human insight and control.
journal: IEEE Computer Graphics & Applicationsfootnotetext: Accepted for publication in IEEE Computer Graphics and Application. DOI: 10.1109/MCG.2025.3607741.
[…]
Agents—software entities that perceive their environment and take goal-directed actions—have been a cornerstone of artificial intelligence research for decades [1]. The emergence of Large Language Models (LLMs) has dramatically accelerated this paradigm, enabling general-purpose agentic AI systems capable of sophisticated planning, reasoning, and autonomous action in complex domains [6] [35]. Given how previous waves of AI innovation have transformed visualization, this agentic turn will inevitably reshape our field. Yet visualization differs fundamentally from other AI application domains through its emphasis on human-centered reasoning, support for open-ended sensemaking, and placement of human analysts at the center of the analytical process. The growing capabilities of agentic systems challenge the human-centered focus of visualization, raising questions about the level of agency, delegation, coordination, and responsibility between humans and AI [6]. What design patterns can guide us in developing agentic systems that leverage agentic capabilities while preserving human insight and analytical control?
To address these questions, we introduce the concept of agentic visualization (AV): interactive visual analysis systems that incorporate autonomous agents(whether AI-driven or rule-based) while maintaining human agency in the analytical process. Rather than starting from scratch, we analyze existing visualization systems, both old and new, with automatic or semi-automatic capabilities—sometimes described as mixed-initiative systems [7]—reinterpreting them as early forms of agentic visualization. While new visualization systems use LLM-based agents, we intentionally focus on prior foundational literature (before the introduction of LLMs), to understand what we can learn from the past to inform our present design patterns for agentic visualization.
By examining these systems through an agentic lens, we identify recurring structures that effectively coordinate automated components and human analysts. We formalize these structures as design patterns for agentic visualization that balance automation and analytical control. Design patterns, originating in architecture [1] and software engineering [5], provide a structured vocabulary for describing reusable solutions to common challenges. This approach aligns with human-centered AI principles that emphasize human values, interpretability, and agency in AI-augmented systems [2, 13].
We validate our proposed design patterns by identifying their presence across diverse published visualization systems. Each pattern is structured with a descriptive name, the problem it addresses, its solution structure, implementation examples, and associated tradeoffs. To demonstrate practical application, we present a novel example scenario that combines multiple patterns to create a dynamic visualization dashboard surveying an active scientific field where agents with different roles, communication, and coordination mechanisms work together to update progressively.
This paper makes the following contributions: (1) The concept of agentic visualization as a framework for designing and analyzing systems that incorporate autonomous agents while preserving human-centered analytical processes; (2) a collection of design patterns for agentic visualization extracted from a systematic analysis of prior foundational literature on automated and semi-automated visualization systems; and (3) a demonstration showing how these patterns can be applied to a novel analytical scenario, including their composition into higher-order solutions.
AGENTS AND AGENTIC VISUALIZATION
Agents are software entities that are capable of perceiving their environment and acting upon it to achieve goals [1]. An agent’s capacity to make their own decisions and take intentional actions towards that goals, is referred to as agency, which lies on a spectrum from simple, reactive behavior to fully autonomous behavior.
We define agentic visualization as the integration of agents with interactive visual analysis systems while balancing computational autonomy with human analytical control. This approach enables agents to independently perform complex analytical tasks while preserving the human analyst’s ability to guide the analytical process, interpret results, and maintain decision-making authority.
Agentic visualization extends and transforms several existing paradigms in visual analytics. Building on mixed-initiative interaction [7] and human-AI collaboration [2], agentic visualization specifically emphasizes agent autonomy beyond reactive assistance. Unlike traditional mixed-initiative systems [15], where humans and agents collaboratively initiate tasks, agentic visualization allows agents to proactively identify opportunities for analysis, suggest approaches, or even pursue analytical threads independently.
Instead of waiting for humans to issue commands and wait for response, agentic visualization challenges this traditional model by establishing a more fluid relationship where both human and agent can initiate analytical activities. This shift prompts questions about how agents should behave when idle: should they remain passive until directed (cold start), or should it proactively explore data and generate insights even without explicit user direction? The answer likely depends on context, task complexity, and user preferences.
Agentic visualization also leverages progressive visual analytics [15], where systems provide incremental visual updates during computationally intensive processes. However, while progressive analytics primarily focuses on making computation visible, agentic visualization extends this by enabling agents to make autonomous decisions about which computational paths to pursue based on intermediate results. This way, agents can prioritize promising analytical directions, abandon unproductive paths, and adapt strategies without continuous human supervision.
To design such an agentic visualization system, we turn to both old and new visualization systems and reframe them through an agentic lens. Through this process, we can identify patterns of effective human-agent collaboration that have emerged organically.
Autonomous agents—software entities that perceive their environment and take goal-directed actions—have been a cornerstone of artificial intelligence research since the field’s inception in the 1950s [1]. The early reactive agents of the 1990s emphasized real-time interaction over complex reasoning, exemplified by Rodney Brooks’ subsumption architecture [3], followed by multi-agent systems and cognitive architectures that aimed to model distributed problem-solving and human cognition respectively [4]. The current resurgence of agent technology, powered by Large Language Models (LLMs), has dramatically accelerated this paradigm, enabling general-purpose agentic AI systems capable of sophisticated planning, reasoning, and autonomous action in complex domains [6]. This represents a significant shift in capabilities but confronts persistent challenges that historical agent systems also faced: poor generalization across domains, scalability issues with complex tasks, inadequate coordination mechanisms, brittle performance in unexpected situations, and ethical concerns. These limitations highlight an inherent tension in agent design that was famously debated by Ben Shneiderman and Pattie Maes at IUI ’97 and CHI ’97, where Shneiderman advocated for direct manipulation interfaces that “affords the user control and predictability,” while Maes argued for delegation to software agents that could learn user preferences and act on their behalf [7]. This debate continues to shape the balance between computational agency and human control in interactive systems. Addressing these challenges requires more than just technical improvements to agent capabilities. As Shah and White argue, successful agent systems need to generate sufficient value for users, provide adaptable personalization, establish trustworthiness, achieve social acceptability, and develop standardization protocols. Aridor and Lange’s work on design patterns for agent systems provides a useful foundation for capturing common solutions to recurring problems in agent design [2]. Their patterns are conceptually divided into three classes: Traveling patterns: Agents moving and navigating in distributed environments; Task patterns: Accomplishing agent objectives and distributing workloads; and Interaction patterns: Governing communication and coordination between agents. Recent work has extended this approach to foundation model-based agents, with researchers developing catalogs of architectural patterns that address challenges such as hallucinations, explainability, and accountability [5].
    Acknowledgments
    This work was supported partly by Villum Investigator grant VL-54492 by Villum Fonden. Any opinions, findings, and conclusions expressed in this material are those of the authors and do not necessarily reflect the views of the funding agency.